THE DIRECT OBSERVATION AND ANALYSIS OF THE JET COOLED ROTATION-TUNNELING TRAN-SITIONS K = 0 ← 0 OF HBr DIMER

\underline{BLAKE A. McELMURRY}, JOHN W. BEVAN, ROBERT R. LUCCHESE, Department of Chemistry, Texas A&M University, College Station, TX 77843-3255; SERGEY P. BELOV, IGOR I. LEONOV, Institute of Applied Physics of Russia Academy of Sciences, Uljanova Str. 46, Nizhny Novgorod 603950, Russia.

The K = 0 ← 0 rotation-tunneling transitions of HBr dimer centered at 450 GHz has been observed for the first time using the TAMU fast scan sub-millimeter wave spectrometer^{a} with coaxially configured pulsed jet. Approximately 1200 completely resolved hyperfine components of the R and P branches with J=0 to 10 in the three isotopomers H^{79}Br:H^{81}Br, H^{79}Br:H^{79}Br and H^{81}Br:H^{81}Br were measured with an accuracy of about 5 kHz. The measured frequencies were fitted to a rms deviation of less than 12 kHz with a standard Hamiltonian which included the rotational (B and D_{J}) and quadrupole terms (χ_{aa}(Br_{1}), χ_{aa}(Br_{2}), D_{χ}(Br_{1}) and D_{χ}(Br_{2})) in the upper and lower states and the tunneling frequency ν_{o}. The accuracy of ν_{o} is improved from MHz to 1 kHz, and all the molecular parameters of the upper tunneling state are now determined with microwave accuracy also. The improved set of parameters for K = 0 ← 0 rotation-tunneling transitions of HBr dimer was extremely useful in the confirmation of the assignment of the observed broken symmetry infrared transitions associated with both ν_{1} and ν_{2} band origins. The parameters from microwave, infrared and the current study will be compared and used to determine structural and other molecular parameters for the dimer. The application of this data for generation of a morphed potential for the hydrogen bromide dimer will also be discussed.

aMcElmurry, B. A.; Lucchese, R. R.; Bevan, J. W.; Leonov, I. I.; Belov, S. P.; Legon, A. C. J Chem Phys 2003, 119, 10687.